In late 1980, the U.S. Commerce Department released new data showing inflation was 15%. At the time, the prime rate of intereSelect the correct yield curve based on these data. A Interest Rate (%) 1 14 134 121 11 101 2 NA 4 6 8 10 12 14 16 18 Years tC 151 Interest Rate (%) 141 134 12 111 101 3 1 2 4 6 8 10 12 14 16 18 Years to Maturity D 15 Interest Rate (%) 14 131 121 111d. Describe the general economic conditions that could lead to an upward-sloping yield curve. I. An upward sloping yield curvPlease provide instructions
Answer #1

1.
=(12%+9%+7%+6%*2)/5=8.00%

2.
=(12%+9%+7%+6%*2)/5+1%=9.00%

3.
=1%+0.1*1%+12%=13.10%

4.
=1%+0.1*2%+(12%+9%)/2=11.70%

5.
=1%+0.1*5%+(12%+9%+7%+6%*2)/5=9.50%

6.
=1%+0.1*10%+(12%+9%+7%+6%*7)/10=9.00%

7.
=1%+0.1*20%+(12%+9%+7%+6%*17)/20=9.50%

8.
Graph B

9.
Option I

10.
Option I